OpenStudy (anonymous):

| x | + 10 = 1 find the solution

OpenStudy (anonymous):

x= 9... Is your answer... :)

OpenStudy (lgbasallote):

not really..

OpenStudy (anonymous):

Then ?

OpenStudy (lgbasallote):

|x| + 10 = 1 subtract 10 from both sides |x| + 10 - 10 = 1 - 10 |x| = -9 now..can you give a number that when you take the absolute value the result is -9?

OpenStudy (anonymous):

no solution

OpenStudy (anonymous):

i got no solution also

OpenStudy (anonymous):

so no solution

OpenStudy (anonymous):

Yeah @lgbasallote..... to be true to the question..

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

If you need a solution desperately.... take 9.... But not true to the question...

OpenStudy (anonymous):

|x-a| = b is a way of saying "the difference between x and a is b". If you rearrange you get |x|=-9. You have no a so a=0, and b = -9. So we have that "the difference between x and 0 is -9". There are 2 numbers that satisfy this, and they are +9 and -9.

OpenStudy (anonymous):

so i should go with negative 9 or 9

OpenStudy (anonymous):

You should go with BOTH -9 and 9.

OpenStudy (anonymous):

thats wrong answer i think it was no solution

OpenStudy (anonymous):

it was!

OpenStudy (anonymous):

flutter

OpenStudy (anonymous):

o we cant cuss on here lmao

OpenStudy (anonymous):

loll

OpenStudy (anonymous):

go ahead try to lol and i fail my damn test so now im pissed !!!!!!

OpenStudy (anonymous):

nooo i might be black but i dont cuss lol and test for what?

OpenStudy (anonymous):

school !!!! and your b;ack with a white guys pic o.O hhah

OpenStudy (anonymous):

yupp! can you say sexy?

OpenStudy (anonymous):

??? hmmmm okay well gooing to sleep now holla

OpenStudy (anonymous):

The reason there is no solution is that it says the modulus is a negative number, namely -9. The absolute value of a number cannot be negative.

OpenStudy (anonymous):

\[\pm 9\]